OpenStudy (anonymous):

ABCD is a parallelogram. Find the value of x. 8? im doing the drawing

OpenStudy (anonymous):

 A=118degrees and b is 5x-2 and d is 7x +4

OpenStudy (anonymous):

x=-3 ?

OpenStudy (anonymous):

nope 5 ,8 ,16 and 24 are my only options

OpenStudy (anonymous):

thanks though were you trying to get the same answer for both if so i did too

jhonyy9 (jhonyy9):

so because this is parallelogramm than we know that A=C and D=B yes ?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

rightagaiin captain

jhonyy9 (jhonyy9):

A=118 so than C=118 total 118+118=236 yes ?

OpenStudy (anonymous):

the answer is 8 because ........

jhonyy9 (jhonyy9):

the sum of all angles degree is 360 degree ,right ?

OpenStudy (anonymous):

360-236

OpenStudy (anonymous):

because its the closest i got to it being equal

OpenStudy (anonymous):

124 /2 or not

OpenStudy (anonymous):

how x cannot be -3? when 5x-2=7x+4

OpenStudy (anonymous):

idk swalah just not an option

OpenStudy (anonymous):

A+d as angles = 180 118+7x+4=180 solve this

jhonyy9 (jhonyy9):

than 5x-2=7x+4 because there we know that D=B,yes ?

OpenStudy (anonymous):

answer never will be 8

OpenStudy (anonymous):

ok swalah , and yes jhonny and red idk what i can multiply to get 58 i mean 7 x ? = 58

jhonyy9 (jhonyy9):

so 360-236=124 so from this result that D+B=124=5x-2+7x+4 yes ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me make sure if u mean that d is an angle so I am not just but I am completely sure

OpenStudy (anonymous):

the options i have never give me an equal

OpenStudy (anonymous):

u have 8

OpenStudy (anonymous):

8 is what i started of with only cus it was closest

OpenStudy (anonymous):

and the angles will be like this b=38 d=60

jhonyy9 (jhonyy9):

so than 124=12x+2 122=12x x=122/12 =61/6

OpenStudy (anonymous):

hmm doesnt make sense though does it

OpenStudy (anonymous):

10.2???